Circular No. 6688

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 D1 (SOHO) C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (see IAUC 6685), the discovery of another faint comet, evidently a Kreutz sungrazer, in SOHO-LASCO C3 coronagraph images by B. McCarty, University of Birmingham, using the automated search software developed by D. Biesecker. Some of the images may have shown a small tail. The comet appeared close to the pylon holding the occulter, so there is some uncertainty in the vignetting. Measurements by St. Cyr and D. Lewis have been reduced by G. V. Williams and are given on MPEC 1997-M10 in detail, the approximate discovery position being as follows:

1996 UT  α_{2000}  δ_{2000}  m_{1}
Feb. 18.150  22^{h}30.^{m}8  −13^{o}41′  5

The comet was observed approaching the sun in p.a. 124^{o} (from the solar north pole) until Feb. 18.59 UT (i.e., over a range of only 24.7 to 17.3 R_{⊙}). Orbital computations by the undersigned, also given in detail on MPEC 1997-M10, indicate perihelion passage on Feb. 19.41.

GRB 970616 A. J. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofísica Es-

pacial y Física Fundamental, Madrid; D. Thompson and K. Birkle, Max Planck Institut für Astronomie, Heidelberg; and J. Greiner, Astrophysikalisches Institut, Potsdam, report, on behalf of a large international collaboration: “Following the detection of GRB 970616 by BATSE and RXTE (IAUC 6683), images in the K′-band (2.1 μm) were obtained with the 3.5-m CAHA telescope (+ OMEGA) at the German-Spanish Calar Alto Observatory on June 21.18, 25.17 and 26.18 UT. A comparison between the first and third frames reveals no variations > 0.5 mag for sources with K′ < 17, either within the fading ASCA x-ray source error circle (IAUC 6687), or in the rest of the RXTE/IPN annulus (IAUC 6687) in a 5′ × 4′ region southeastward from this source. The optical variable reported by Galama et al. is present on all three images but does not show any obvious variation: preliminary K′ magnitude estimates are 17.5 ± 0.4, 17.5 ± 0.3 and 17.4 ± 0.4, respectively. This suggests a different origin for the observed optical fading. With the estimated R ∼ 24 on June 24.61 (IAUC 6687), the R − K′ color on that date was ∼ 6.5, placing it on the red side of the distribution of all objects at K′ ∼ 18.”

1997 June 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden